x+(-2.8)+9.2=0 what is x
Answer by haileytucki(390)

x+(-2.8)+9.2=0
Remove the parentheses around the expression -2.8.
x-2.8+9.2=0
Add 9.2 to -2.8 to get 6.4.
x+6.4=0
Since 6.4 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 6.4 from both sides.
x=-6.4